# SegDINO3D: 3D Instance Segmentation Empowered by Both Image-Level and Object-Level 2D Features

### v1

@@ Abstract
In this paper, we present SegDINO3D, a novel Transformer encoder-decoder framework for 3D instance segmentation. As 3D training data is generally not as sufficient as 2D training images, SegDINO3D is designed to fully leverage 2D representation from a pre-trained 2D detection model, including both image-level and object-level features, for improving 3D representation. SegDINO3D takes both a point cloud and its associated 2D images as input. In the encoder stage, it first enriches each 3D point by retrieving 2D image features from its corresponding image views and then leverages a 3D encoder for 3D context fusion. In the decoder stage, it formulates 3D object queries as 3D anchor boxes and performs cross-attention from 3D queries to 2D object queries obtained from 2D images using the 2D detection model. These 2D object queries serve as a compact object-level representation of 2D images, effectively avoiding the challenge of keeping thousands of image feature maps in the memory while faithfully preserving the knowledge of the pre-trained 2D model. The introducing of 3D box queries also enables the model to modulate cross-attention using the predicted boxes for more precise querying. SegDINO3D achieves the state-of-the-art performance on the ScanNetV2 and ScanNet200 3D instance segmentation benchmarks. Notably, on the challenging ScanNet200 dataset, SegDINO3D significantly outperforms prior methods by +8.7 and +6.8 mAP on the validation and hidden test sets, respectively, demonstrating its superiority.
footnotetext: * Equal contribution. Work done during an internship at IDEA Research. † Corresponding author.
@@ 1 Introduction
In summary, our contributions are three-fold:
@@
  • SegDINO3D achieves the state-of-the-art performance on the 3D instance segmentation benchmarks ScanNet200 and ScanNetV2. Notably, SegDINO3D surpasses prior methods by +8.7 and +6.8 mAP on the validation and hidden test sets of ScanNet200, respectively.
@@ 2 Related Works
Recently, Transformer-based methods (Sun et al. 2023; Al Khatib et al. 2023; Wang et al. 2024a) have emerged and started to dominate this field. Similar to DETR-like architectures in 2D perception, they directly predict segmentation results and offer advantages in both accuracy and inference speed. Mask3D (Schult et al. 2023) first proposes a Transformer architecture for 3D instance segmentation. QueryFormer (Lu et al. 2023) and MAFT (Lai et al. 2023) further improved query initialization and refinement. OneFormer3D (Kolodiazhnyi et al. 2024a) unifies semantic, instance, and panoptic segmentation within a single framework. These methods operate directly on 3D point clouds, while ignoring the use of 2D images, whereas ODIN (Jain et al. 2024) leverages RGB-D images and employs a unified model to perform both 2D and 3D segmentation.
@@ 3.3 Box-Modulated Cross-Attention with 3D Superpoint
where $SimCont$ indicates the content similarity between query and key, $\mathbf{A}\in\mathbb{R}^{M\times S}$ is the attention weights. However, the attention mask $\mathbf{M}_{\infty}$ solely relies on the similarity between the content embeddings of the 3D object queries and the superpoint features, without considering any spatial constraints. This may result in high similarity scores between different instances of the same category that are spatially far apart, potentially leading to confusion. Therefore, it is crucial to introduce boxes to constrain the cross-attention.
@@ 4.2 Comparison with Prior Work
We compare our method with previous approaches on the ScanNet200 validation and hidden test set. Our method offers two versions that utilize either Grounding DINO or DINO-X as the 2D model to provide features. In the following discussions, we refer to the DINO-X-based version for comparison. As shown in Table 4, on the validation set, SegDINO3D outperforms all existing methods by a large margin in all sub-splits, achieving a +8.7 mAP higher mAP than the previous best method. Especially, SegDINO3D’s performance on the tail-split even surpasses previous methods on the common-split (36.2 mAP vs. 31.6 mAP). This indicates that the incorporation of 2D features significantly alleviates the issue of insufficient 3D training data for long-tailed categories. Meanwhile, on the hidden test set, SegDINO3D also shows significant improvement, establishing a new state of the art. These consistent and substantial improvements demonstrate the superiority of SegDINO3D.
Moreover, on the ScanNetv2 validation set, SegDINO3D outperforms the baseline by +4.7 mAP and surpasses the previous state-of-the-art method by +1.5 mAP. For detailed comparisons, please refer to our technical appendix.
@@
Table 2: Ablation on our key components. Img. F. denotes the use of 2D image-level features in encoder, Obj. F. refers to the incorporation of 2D object-level features in decoder, and Mod. indicates the use of box modulated cross-attention. denotes using predicted boxes to filter masks.
@@ Appendix A Technical Appendix
In the technical appendix, we provide more quantitative (Sec. A.1) and qualitative (Sec. A.4) comparisons with previous methods to further demonstrate the superiority of our approach. Furthermore, we also conduct additional ablation studies based on the Grounding DINO to validate the effectiveness and generalization ability of our method (Sec. A.2). Finally, we provide more implementation details to facilitate reproducibility (Sec. A.3).
A.1 More Comparisons with Previous Methods
We compare SegDINO3D with previous methods on the ScanNetv2 validation set. As shown in Table 7, our method outperforms all existing methods across all metrics. Specifically, it achieves a 4.7 mAP improvement over our baseline (OneFormer3D) and surpasses the previous SoTA method by 1.5 mAP, demonstrating the superiority of our method.
Compared with the improvement achieved on the more challenging ScanNet200, the improvement on ScanNetv2 is relatively smaller (+10.0 mAP v.s. +4.7 mAP). We attribute this to two main reasons: 1) 2D representations provide more discriminative semantic features, leading to more advantage in the fine-grained 200-class setting of ScanNet200; 2) ScanNetv2 only has 20 classes, and is relatively simpler. The performance of existing methods on ScanNetv2 has begun to saturate, leaving less room for substantial improvements.
Method mAP mAP50 mAP25
SSTNet (2021) 49.4 64.3 74.0
SoftGroup (2022) 46.0 67.6 78.9
Mask3D (2023) 55.2 73.7 83.5
SPFormer (2023) 56.3 73.9 82.9
QueryFormer (2023) 56.5 74.2 83.3
MAFT (2023) 59.9 76.5 -
OneFormer3D (2024a) 59.3 78.1 86.4
TD3D (2024b) 47.3 71.2 81.9
ODIN (2024) 50.0 71.0 83.6
Spherical Mask (2024) 62.3 79.9 88.2
Relation3D (2025) 62.5 80.2 87.0
Ours-DINO-X 64.0 81.5 88.9
Table 7: Comparison of SegDINO3D with prior methods on validation set of ScanNetv2.
@@ A.2 More Ablation Studies
To validate the effectiveness, we conduct an ablation study. Considering practical feasibility, we only compare with the way that mapping 2D image features to discrete 3D points. More specifically, we project the original 2D feature maps onto discrete 3D points and fuse them with the point cloud features $\mathbf{F}^{3D}$ obtained from the 3D encoder using an MLP. As shown in Table 8, compared with the use of discrete 2D features, the use of the 2D object queries shows a significant superiority (+4.1 mAP).
@@
Table 9: Ablation on Dropout of Object-Level Qeuries.
Key Components with Grounding DINO. We use Grounding DINO (Liu et al. 2024) as the 2D DETR-based model to conduct the same ablation studies as in Table  2, to validate the effectiveness and generalization ability of our proposed key components. However, different from the setting in Sec. 4.3 of the main text, Grounding DINO is not as powerful as DINO-X (it is also trained on a smaller amount of 2D data), so we require a longer training schedule to achieve convergence. Specifically, we train for 108 epochs. As shown in Table 10, compared to the results in Table 2, the conclusions remain consistent.
@@
Table 10: Ablation on our key components with Grounding DINO as 2D model. Img. F. denotes the use of 2D image-level features in encoder, Obj. F. refers to the incorporation of 2D object-level features in decoder, and Mod. indicates the use of box modulated cross-attention. denotes using predicted boxes to filter masks.

### v2

@@ Abstract
In this paper, we present SegDINO3D, a novel Transformer encoder-decoder framework for 3D instance segmentation. As 3D training data is generally not as sufficient as 2D training images, SegDINO3D is designed to fully leverage 2D representation from a pre-trained 2D detection model, including both image-level and object-level features, for improving 3D representation. SegDINO3D takes both a point cloud and its associated 2D images as input. In the encoder stage, it first enriches each 3D point by retrieving 2D image features from its corresponding image views and then leverages a 3D encoder for 3D context fusion. In the decoder stage, it formulates 3D object queries as 3D anchor boxes and performs cross-attention from 3D queries to 2D object queries obtained from 2D images using the 2D detection model. These 2D object queries serve as a compact object-level representation of 2D images, effectively avoiding the challenge of keeping thousands of image feature maps in the memory while faithfully preserving the knowledge of the pre-trained 2D model. The introducing of 3D box queries also enables the model to modulate cross-attention using the predicted boxes for more precise querying. SegDINO3D achieves the state-of-the-art performance on the ScanNetV2 and ScanNet200 3D instance segmentation benchmarks. Notably, on the challenging ScanNet200 dataset, SegDINO3D significantly outperforms prior methods by +8.6 and +6.8 mAP on the validation and hidden test sets, respectively, demonstrating its superiority.
Codehttps://github.com/IDEA-Research/SegDINO3D
@@ 1 Introduction
In summary, our contributions are threefold:
@@
  • SegDINO3D achieves the state-of-the-art performance on the 3D instance segmentation benchmarks ScanNet200 and ScanNetV2. Notably, SegDINO3D surpasses prior methods by +8.6 and +6.8 mAP on the validation and hidden test sets of ScanNet200, respectively.
@@ 2 Related Works
Recently, Transformer-based methods (Sun et al. 2023; Al Khatib et al. 2023; Wang et al. 2024a) have emerged and started to dominate this field. Similar to DETR-like architectures in 2D perception, they directly predict segmentation results and offer advantages in both accuracy and inference speed. Mask3D (Schult et al. 2023) first proposes a Transformer architecture for 3D instance segmentation. QueryFormer (Lu et al. 2023) and MAFT (Lai et al. 2023) further improved query initialization and refinement. OneFormer3D (Kolodiazhnyi et al. 2024a) unifies semantic, instance, and panoptic segmentation within a single framework. Relation3D (Lu and Deng 2025) explicitly models relationships through a geometric-biased self-attention for queries and a contrastive learning scheme for scene features. These methods operate directly on 3D point clouds, while ignoring the use of 2D images, whereas ODIN (Jain et al. 2024) leverages RGB-D images and employs a unified model to perform both 2D and 3D segmentation.
@@ 3.3 Box-Modulated Cross-Attention with 3D Superpoint
where $SimCont$ indicates the content similarity between query and key, $\mathbf{A}\in\mathbb{R}^{M\times S}$ is the attention weights. However, the attention mask $\mathbf{M}_{\infty}$ solely relies on the similarity between the content embeddings of the 3D object queries and the superpoint features, without considering any spatial constraints. This may result in high similarity scores between different instances of the same category that are spatially far apart, potentially leading to confusion. Some methods (Lai et al. 2023; Lu and Deng 2025) attempt to incorporate positional embeddings to provide spatial constraints. However, it is still isotropic and insufficient for capturing object scale variations, which are prevalent in 3D environments. Therefore, we further propose a box-modulated cross-attention mechanism that breaks the isotropy of positional embeddings, enabling more precise attention control.
@@ 4.2 Comparison with Prior Work
We compare our method with previous approaches on the ScanNet200 validation and hidden test set. Our method offers two versions that utilize either Grounding DINO or DINO-X as the 2D model to provide features. In the following discussions, we refer to the DINO-X-based version for comparison. As shown in Table 3.4, on the validation set, SegDINO3D outperforms all existing methods by a large margin in all sub-splits, achieving a +8.6 mAP higher mAP than the previous best method. Especially, SegDINO3D’s performance on the tail-split even surpasses previous methods on the common-split (36.2 mAP vs. 31.6 mAP). This indicates that the incorporation of 2D features significantly alleviates the issue of insufficient 3D training data for long-tailed categories. Meanwhile, on the hidden test set, SegDINO3D also shows significant improvement, establishing a new state of the art. These consistent and substantial improvements demonstrate the superiority of SegDINO3D.
Method mAP mAP50 mAP25
SSTNet (2021) 49.4 64.3 74.0
SoftGroup (2022) 46.0 67.6 78.9
Mask3D (2023) 55.2 73.7 83.5
SPFormer (2023) 56.3 73.9 82.9
QueryFormer (2023) 56.5 74.2 83.3
MAFT (2023) 59.9 76.5 -
OneFormer3D (2024a) 59.3 78.1 86.4
TD3D (2024b) 47.3 71.2 81.9
ODIN (2024) 50.0 71.0 83.6
Spherical Mask (2024) 62.3 79.9 88.2
Relation3D (2025) 62.5 80.2 87.0
Ours-DINO-X 64.0 81.5 88.9
Table 2: Comparison of SegDINO3D with prior methods on the validation set of ScanNetv2.
Moreover, we compare SegDINO3D with prior methods on the ScanNetv2 validation set. As shown in Table 2, our method outperforms all existing methods across all metrics. Specifically, it outperforms our baseline by +4.7 mAP and surpasses the previous state-of-the-art method by +1.5 mAP. Compared with the improvement achieved on the more challenging ScanNet200, the improvement on ScanNetv2 is relatively smaller. We attribute this to the fact that 2D representations provide more discriminative semantic features, which are particularly advantageous under the fine-grained 200-class setting of ScanNet200.
@@
Table 3: Ablation on our key components. Img. F. denotes the use of 2D image-level features in the encoder, Obj. F. refers to the incorporation of 2D object-level features in the decoder, and Mod. indicates the use of box modulated cross-attention. denotes using predicted boxes to filter masks.
@@ Appendix A Technical Appendix
In the technical appendix, we first provide an efficiency analysis of SegDINO3D (Sec. A.1). Moreover, we conduct additional ablation studies based on the Grounding DINO to validate the effectiveness and generalization ability of our method (Sec. A.2). Furthermore, we provide more implementation details to facilitate reproducibility (Sec. A.3). Finally, we provide more qualitative (Sec. A.4) comparisons with previous methods to further demonstrate the superiority of our approach.
A.1 Efficiency Analysis
We evaluate the efficiency of our model on the ScanNet200 validation scenes and report the averaged results. We divide SegDINO3D into two stages, namely Feature Preparation Stage and Model Inference Stage, and provide a detailed analysis of each.
Feature Preparation Stage. We first employ DINO-X (Ren et al. 2024) to extract features from scene images. For each scene, we sample input frames at intervals of 5. When using RGB images at their original resolution with $width=1536$, the average processing time per scene is 94.3 s. By reducing the input resolution to $width=768$, the average time decreases to only 22.3 s, while performance remains nearly unchanged (a drop of merely -0.5 mAP). Subsequently, we apply the Nearest View Sampling strategy described in Sec. 3.2 to extract appropriate features for each 3D point. This process involves simple matrix operations and takes only 0.8 ms per scene, which is negligible.
Model Inference Stage. Our model introduces only a marginal computational overhead compared to the baseline (OneFormer3D). Specifically, during inference on an RTX 3090 GPU, the per-scene processing time increases slightly from 310 ms to 350 ms, indicating that the proposed method maintains high efficiency while achieving superior performance. Similarly, the peak GPU memory reserved shows a minimal rise from 15.0 GB to 15.5 GB, demonstrating that the method remains memory-efficient. The detailed comparison is summarized in Table 8.
Metric Baseline Ours
Runtime per scene (ms) 310 350
Peak GPU memory reserved (GB) 15.0 15.5
Table 8: Inference efficiency comparison with the baseline.
In summary, in the offline scenarios, the runtime and memory consumption of SegDINO3D are both practical.
@@ A.2 More Ablation Studies
To validate the effectiveness, we conduct an ablation study. Considering practical feasibility, we only compare with the way that maps 2D image features to discrete 3D points. More specifically, we project the original 2D feature maps onto discrete 3D points and fuse them with the point cloud features $\mathbf{F}^{3D}$ obtained from the 3D encoder using an MLP. As shown in Table 9, compared with the use of discrete 2D features, the use of the 2D object queries shows a significant superiority (+4.1 mAP).
@@
Table 10: Ablation on Dropout of Object-Level Queries.
Key Components with Grounding DINO. We use Grounding DINO (Liu et al. 2024) as the 2D DETR-based model to conduct the same ablation studies as in Table 3, to validate the effectiveness and generalization ability of our proposed key components. However, different from the setting in Sec. 4.3 of the main text, Grounding DINO is not as powerful as DINO-X (it is also trained on a smaller amount of 2D data), so we require a longer training schedule to achieve convergence. Specifically, we train for 108 epochs. As shown in Table 11, compared to the results in Table 3, the conclusions remain consistent.
@@
Table 11: Ablation on our key components with Grounding DINO as 2D model. Img. F. denotes the use of 2D image-level features in the encoder, Obj. F. refers to the incorporation of 2D object-level features in the decoder, and Mod. indicates the use of box modulated cross-attention. denotes using predicted boxes to filter masks.